# Electromagnetic Inverse Scattering from a Single Transmitter

### v5

@@ 4.3 Training
5 Experiments
5.1 Setup
Datasets. We train and test our method on standard benchmarks used for EISP following previous work (Wei and Chen, 2019; Song et al., 2021; Liu et al., 2022). To enhance training efficiency and model robustness, datasets that share identical transmitter and receiver configurations are combined into a unified training set. 1) Synthetic Circular-cylinder dataset (Circular(Luo et al., 2024) is synthetically generated comprising images of cylinders with random relative radius, number, and location and permittivity. 2) Synthetic MNIST dataset (MNIST(Deng, 2012) contains grayscale images of handwritten digits. For the two synthetic datasets, we follow previous work(Luo et al., 2024; Deng, 2012), we evaluate two levels of noise: 5% and 30%. 3) Real-world Institut Fresnel’s database (IF(Geffrin et al., 2005) contains three different dielectric scenarios, namely FoamDielExt, FoamDielInt, and FoamTwinDiel. 4) Synthetic 3D MNIST dataset (3D MNIST(de la Iglesia Castro, ) contains 3D data of handwritten digits. 5) 3D ShapeNet dataset (3D ShapeNet(Wu et al., 2015) contains 3D data of various shapes. For more details about datasets, please refer to Appx.˜C.
@@ 5.1 Setup
  Number of Transmitters: $N=16$ $N=8/18$
BP (2005) 0.177 0.719 16.43 0.178 0.716 16.38 0.052 0.905 27.41 0.053 0.904 27.42 0.190 0.779 16.19
2-fold SOM (2009) 0.154 0.757 20.93 0.156 0.738 20.84 0.031 0.917 32.23 0.038 0.889 30.63 - - -
Gs SOM (2009) 0.072 0.923 28.31 0.081 0.901 27.13 0.023 0.946 35.40 0.024 0.937 34.89 0.184 0.790 17.00
BPS (2019) 0.093 0.909 25.00 0.105 0.891 23.90 0.027 0.963 33.00 0.029 0.956 32.42 0.310 0.664 17.05
PGAN (2021) 0.084 0.916 25.80 0.091 0.910 25.31 0.026 0.966 35.56 0.032 0.947 33.91 0.121 0.926 24.78
Physics-Net (2022) 0.075 0.932 26.17 0.093 0.906 24.58 0.027 0.934 32.72 0.030 0.927 32.08 0.170 0.788 18.48
Img-Interiors (2024) 0.200 0.863 26.41 0.336 0.760 19.01 0.036 0.947 35.05 0.047 0.932 32.62 0.153 0.837 23.26
@@
  Number of Transmitters: $N=1$
BP (2005) 0.194 0.698 15.40 0.194 0.696 15.40 0.065 0.892 25.30 0.065 0.892 25.30 0.199 0.770 16.29
2-fold SOM (2009) 0.432 0.556 12.49 0.828 0.382 9.45 0.060 0.859 26.63 0.157 0.639 20.07 - - -
Gs SOM (2009) 0.460 0.598 15.31 0.404 0.557 14.91 0.046 0.888 29.62 0.051 0.862 28.77 0.192 0.779 16.66
BPS (2019) 0.189 0.774 18.75 0.205 0.744 17.97 0.045 0.891 29.29 0.055 0.862 27.68 0.348 0.669 16.18
PGAN (2021) 0.133 0.867 21.69 0.153 0.830 20.41 0.033 0.932 32.02 0.040 0.914 29.94 0.248 0.680 16.85
PhysicsNet (2022) 0.137 0.798 19.98 0.152 0.783 19.38 0.055 0.887 26.60 0.056 0.890 26.48 0.175 0.771 17.45
Img-Interiors (2024) 0.305 0.604 16.06 0.467 0.484 12.47 0.096 0.855 26.19 0.153 0.806 20.90 0.305 0.705 17.34
@@
Baselines and Metrics. We maintain the same settings as in previous studies (Wei and Chen, 2019; Song et al., 2021; Sanghvi et al., 2019) to ensure a fair comparison. We compare our method with three traditional methods and four deep learning-based approaches: 1) BP (Belkebir et al., 2005): A traditional non-iterative inversion algorithm. 2) 2-fold SOM (Zhong and Chen, 2009): A traditional iterative minimization scheme by using SVD decomposition. 3) Gs SOM (Chen, 2009): A traditional subspace-based optimization method by decomposing the operator of Green’s function. 4) BPS (Wei and Chen, 2019): A CNN-based image translation method with an initial guess from the BP algorithm. 5) Physics-Net (Liu et al., 2022): A CNN-based approach that incorporates physical phenomena during training. 6) PGAN (Song et al., 2021): A CNN-based approach using a generative adversarial network. 7) Img-Interiors (Liu et al., 2022): An implicit approach optimized by forward calculation. Following previous work (Liu et al., 2022), we evaluate the quantitative performance of our method using PSNR (Wang and Bovik, 2009), SSIM (Wang et al., 2004), and Relative Root-Mean-Square Error (MSE(Song et al., 2021).
@@ 5.2.1 Multiple Transmitter Evaluation
We begin by comparing our method against prior approaches under the multiple-transmitter setting, using both synthetic and real datasets for comprehensive evaluation. As shown in the upper part of Tab.˜1, our method achieves comparable or superior performance to the SOTA in most cases, demonstrating how our end-to-end training framework successfully leverage the data prior across diverse data domains.
In addition, we present a qualitative comparison, as shown in Fig.˜4. Traditional methods such as BP(Belkebir et al., 2005), Gs SOM (Chen, 2009), and 2-fold SOM (Chen, 2009) are only capable of recovering the coarse shape of the scatterer. BPS (Chen, 2018) produces sharp edges, but the reconstructed shapes are often inaccurate. PGAN (Song et al., 2021) achieves accurate shape recovery, yet introduces noticeable background artifacts. Img-Interiors (Luo et al., 2024) can generate high-quality reconstructions, but occasionally fails due to local optima, as it is based on an iterative optimization process (see the last row). In contrast, our method produces accurate and clean reconstructions across all cases, demonstrating both visual fidelity and robustness.
@@ 5.2.2 Single Transmitter Evaluation
Furthermore, we investigate a highly challenging and practically important setting that has been largely underexplored in previous work: performing EISP with a minimal number of transmitters. Specifically, we consider the most extreme case, using only a single transmitter. As shown in the lower part of Tab.˜1, our method significantly outperforms all previous approaches across all datasets and noise levels. This remarkable performance under such constrained conditions underscores the efficacy of our end-to-end training framework, which successfully encodes and leverages rich data priors to achieve state-of-the-art results across diverse domains.
To better understand this phenomenon, we present qualitative comparisons in Fig.˜5. Traditional methods such as BP(Belkebir et al., 2005), Gs SOM (Chen, 2009), and 2-fold SOM (Chen, 2009) produce only blurry reconstructions. Deep learning-based methods like BPS (Chen, 2018), Physics-Net (Liu et al., 2022), and PGAN (Song et al., 2021) tend to “hallucinate” the digit, resulting in wrong shape on the MNIST dataset. Img-Interiors (Luo et al., 2024) fails to capture the fundamental morphology of the scatterer, resulting in structurally inaccurate representations that deviate significantly from the ground truth. Among all the methods, only ours can still produce reasonably accurate reconstructions of the relative permittivity under such an extreme condition.
@@ 5.3 Ablation Study
Noise Robustness. To simulate real-world sensor noise and related perturbations, we evaluate the robustness of the models by adding noise to the scattered field. Moving beyond simple binary testing, we systematically assess the model’s performance across multiple noise levels ranging from 5% to 30% to examine its behavior in various noisy environments. The quantitative results presented in Tab.˜3 demonstrate that our model exhibits smooth and gradual performance degradation as the noise level increases, maintaining excellent reconstruction capability even under strong noise interference as high as 30%. Qualitative visualizations in Fig.˜6 show that most baseline methods exhibit noticeable artifacts or even complete failure under severe noise conditions, while our method remains robust and preserves the essential structure of the target.
@@
 Data Size MNIST (5%) MNIST (30%)
@@ 5.4 Reconstruction on 3D data
Setup and Metrics. Our method can be naturally extended to 3D scenarios. Following previous work (Luo et al., 2024), we employ the Synthetic 3D MNIST (de la Iglesia Castro, ) and extend to 3D ShapeNet (Wu et al., 2015) for training and testing. For evaluation, we adopt 3D versions of the MSE (Song et al., 2021) and Intersection over Union (IoU) as our metrics. Further details on the datasets are provided in Appx.˜C.
Results. We evaluate our method and Img-Interiors under limited-transmitter settings. Quantitative results demonstrate the superiority of our approach: on 3D MNIST, our method achieves an MSE of 0.120 and IoU of 0.769 with $N=1$ transmitter, significantly outperforming Img-Interiors which obtains an MSE of 0.372 and IoU of 0.094 under the same conditions. With $N=6$ transmitters, our results further improve to MSE of 0.094 and IoU of 0.834. For the more complex 3D ShapeNet dataset under $N=1$ configuration, our method obtains an MSE of 0.064 and IoU of 0.762, showcasing its generalization capability to diverse 3D structures. Fig.˜1 provides a comprehensive comparison of reconstruction quality and runtime between the two methods for both $N=1$ and $N=6$ configurations. Fig.˜7 and Fig.˜8 provide visual comparisons of additional 3D reconstruction results on the 3DMNIST and 3DShapeNet datasets. These results show that our method maintains robustness and generalizes effectively across geometrically complex 3D structures, representing significant progress towards practical applications.
@@ 6 Conclusion
Limitations. While our method effectively handles sparse transmitter settings, it cannot deal with different locations of receivers or transmitters, which remains an important direction for future work.
@@ Appendix B Physical Model of EISP
In EISP, the scatterer is placed within the region of interest $\mathcal{D}$. The transmitters are placed around $\mathcal{D}$, emitting incident electromagnetic field ${E}^{\text{i}}$. The incident field then interacts with the scatterer, exciting the induced current $J$. The induced current can act as secondary radiation sources, emitting scattered field $E^{\text{s}}$. In fact, for a point in the scatterer, it cannot distinguish the incident field and the scattered field at this point. So, it interacts with the sum of the incident field and the scattered field, aka the total field $E^{\text{t}}$. The total field can be described by Lippmann-Schwinger equation (Colton and Kress, 2013) as follows:
@@
Since digital analysis only applies to discrete variables (Mitra, 2001; Taflove et al., 2005), we discretize equations Eq.˜4, Eq.˜5, and Eq.˜6. The region of interest $\mathcal{D}$ is discretized into $M\times M$ square subunits, and we use the method of moment (Peterson et al., 1998) to obtain the discrete scattered field $\mathbf{E}^{\text{s}}$ [^1] We use the bold letters to represent discrete variables.. The discrete version of Eq.˜4 is as follows:
@@ C.1 Datasets
The datasets utilized in this work are described in detail below. To enhance model robustness and training efficiency, datasets that share identical measurement settings are pooled to form consolidated training sets. We strictly follow the established rules in EISP (Wei and Chen, 2019; Song et al., 2021; Liu et al., 2022) to set the dataset. We train and test our model on standard benchmarks used for EISP.
@@
1) Circular (Luo et al., 2024) is synthetically generated comprising images of cylinders with random relative radius, number, location, and permittivity between $1$ and $1.5$. $10k$ images are generated for training purposes, and $1.2k$ images for testing.
@@
2) MNIST (Deng, 2012) contains grayscale images of handwritten digits. Similar to previous settings (Wei and Chen, 2019; Zhou et al., 2022), we use them to synthesize scatterers with relative permittivity values between $2$ and $2.5$ according to their corresponding pixel values. The entire MNIST training set containing $60k$ images is used for training purposes, while $1.2k$ images from the MNIST test set are randomly selected for testing.
@@
Figure 9: Positions of transmitters and receivers for 2D data. (a) For Circular dataset and MNIST dataset, we set up $N=16$ transmitters and $N_{r}=32$ receivers. All transmitters and receivers are equally placed. (b) For IF dataset, we set up $N=8$ or $18$ transmitters and $N_{r}=241$ receivers. The transmitters are equally placed (not shown in the figure), and the locations of receivers are determined by the transmitter. For all datasets, the green square represents the region of interest $\mathcal{D}$.
@@
Following previous work (Chen, 2018; Zhou et al., 2022) to generate the above two synthetic datasets, we set operating frequency $f=400$ MHz. The region of interest is a square with the size of $2m\times 2m$. We use $16$ transmitters and $32$ receivers equally placed on a circle with a radius of $3m$ ($N_{r}=32$ and $N=16$). The schematic diagram of the locations of the transmitters and receivers around the region of interest is shown in LABEL:fig:schematic_synthetic. The data are generated numerically using the method of moments (Peterson et al., 1998) with a $224\times 224$ grid mesh to avoid inverse crime (Colton et al., 1998). To simulate the noise in actual measurement, we add a $5\%$ level of noise to the scattered field $\mathbf{E}^{\text{s}}$ for regular setting.
@@
3) Real-world IF dataset (Geffrin et al., 2005) contains three different dielectric scenarios, namely FoamDielExt, FoamDielInt, and FoamTwinDiel. $N=8$ for FoamDielExt and FoamDielInt, $N=18$ for FoamTwinDiel, and $N_{r}=241$ for all the cases. The region of interest is a square with the size of $0.15m\times 0.15m$. Transmitters and receivers are placed on a circle with a radius of $1.67m$. The transmitters are placed equally, and the locations of receivers vary for each transmitter. There is no receiver at any position closer than $60^{\circ}$ from the transmitter, and 241 receivers are placed from $60^{\circ}$ to $300^{\circ}$ with a step of $1^{\circ}$ from the location of the transmitter. The schematic diagram of the locations of the transmitters and receivers around the region of interest is shown in LABEL:fig:schematic_IF. In the real measurement, there is only one transmitter at a fixed location, and the scatterer rotates to simulate the transmitter to be in different directions. There is a movable receiver sequentially measures the scattered field at 241 different locations. After the measurement, the scatterer rotates by a certain angle for next measurement. The angle is $45^{\circ}$ for FoamDielExt and FoamDielInt because $N_{r}=8$, and $20^{\circ}$ for FoamTwinDiel because $N_{r}=18$. As for operating frequency, all cases are measured under many different frequencies, and we take the frequency $f=5$ GHz. We evaluate these three scenarios for testing, and use the same settings to synthetically generate $10k$ images of cylinders with random number and location for training purposes.
@@
4) Synthetic 3D MNIST (de la Iglesia Castro, ) contains 3D images of handwritten digits. 5) Synthetic 3D ShapeNet (Wu et al., 2015) contains 3D images of various shapes. We use these two datasets to synthesize scatterers with relative permittivity value of $2$. $N_{r}=160$ and $N=40$, and the operating frequency $f=400$ MHz. The region of interest is a cube with the size of $2m\times 2m\times 2m$. The transmitters and receivers are placed at a sphere with the radius of $3m$. For the positions of transmitters, the azimuthal angle ranges from $0^{\circ}$ to $315^{\circ}$ with the step of $45^{\circ}$, and the polar angle ranges from $30^{\circ}$ to $150^{\circ}$ with the step of $30^{\circ}$. For the positions of receivers, the azimuthal angle ranges from $0^{\circ}$ to $348.75^{\circ}$ with the step of $11.25^{\circ}$, and the polar angle ranges from $30^{\circ}$ to $150^{\circ}$ with the step of $30^{\circ}$. The schematic diagram of the locations of the transmitters and receivers around the region of interest is shown in Fig.˜10. The entire 3D MNIST dataset is used, including $5k$ images for training purposes and $1k$ images for testing. For 3D ShapeNet, we take $5k$ images from $5$ different categories for training purposes and $500$ images for testing.

### v6

@@ 4.3 Training
[…]
BP [4] 0.177 0.719 16.43 0.178 0.716 16.38 0.052 0.905 27.41 0.053 0.904 27.42 0.190 0.779 16.19
2-fold SOM [54] 0.154 0.757 20.93 0.156 0.738 20.84 0.031 0.917 32.23 0.038 0.889 30.63 - - -
Gs SOM [6] 0.072 0.923 28.31 0.081 0.901 27.13 0.023 0.946 35.40 0.024 0.937 34.89 0.184 0.790 17.00
BPS [7, 45] 0.093 0.909 25.00 0.105 0.891 23.90 0.027 0.963 33.00 0.029 0.956 32.42 0.310 0.664 17.05
PGAN [39] 0.084 0.916 25.80 0.091 0.910 25.31 0.026 0.966 35.56 0.032 0.947 33.91 0.121 0.926 24.78
Physics-Net [26] 0.075 0.932 26.17 0.093 0.906 24.58 0.027 0.934 32.72 0.030 0.927 32.08 0.170 0.788 18.48
Two-Step [51] 0.111 0.835 23.17 0.111 0.835 22.91 0.059 0.830 25.92 0.091 0.848 23.94 - - -
Img-Interiors [27] 0.200 0.863 26.41 0.336 0.760 19.01 0.036 0.947 35.05 0.047 0.932 32.62 0.153 0.837 23.26
@@
[…]
BP [4] 0.194 0.698 15.40 0.194 0.696 15.40 0.065 0.892 25.30 0.065 0.892 25.30 0.199 0.770 16.29
2-fold SOM [54] 0.432 0.556 12.49 0.828 0.382 9.45 0.060 0.859 26.63 0.157 0.639 20.07 - - -
Gs SOM [6] 0.460 0.598 15.31 0.404 0.557 14.91 0.046 0.888 29.62 0.051 0.862 28.77 0.192 0.779 16.66
BPS [7, 45] 0.189 0.774 18.75 0.205 0.744 17.97 0.045 0.891 29.29 0.055 0.862 27.68 0.348 0.669 16.18
PGAN [39] 0.133 0.867 21.69 0.153 0.830 20.41 0.033 0.932 32.02 0.040 0.914 29.94 0.248 0.680 16.85
PhysicsNet [26] 0.137 0.798 19.98 0.152 0.783 19.38 0.055 0.887 26.60 0.056 0.890 26.48 0.175 0.771 17.45
Two-Step [51] 0.117 0.845 23.34 0.203 0.656 17.62 0.145 0.673 19.45 0.174 0.652 18.41 - - -
Img-Interiors [27] 0.305 0.604 16.06 0.467 0.484 12.47 0.096 0.855 26.19 0.153 0.806 20.90 0.305 0.705 17.34
@@
5 Experiments
5.1 Setup
Datasets. We train and test our method on standard benchmarks used for EISP following previous work [45, 39, 26]. Datasets that share identical transmitter and receiver configurations are combined into a unified training set. 1) Synthetic Circular [27] is synthetically generated comprising images of cylinders with random relative radius, number, location, and permittivity. 2) Synthetic MNIST [12] contains grayscale images of handwritten digits. For the two synthetic datasets, following previous work[27, 12], we evaluate two levels of noise: 5% and 30%, and the number of receivers $N_{r}=32$. 3) Real-world Institut Fresnel’s database (IF[15] contains three different dielectric scenarios, namely FoamDielExt, FoamDielInt, and FoamTwinDiel, where $N_{r}=241$. 4) Synthetic 3D MNIST dataset (3D MNIST[11] contains 3D data of handwritten digits. 5) 3D ShapeNet dataset (3D ShapeNet[46] contains 3D data of various shapes. $N_{r}=160$ for the two 3D datasets. For more details about datasets, please refer to Appx.˜C.
Baselines and Metrics. For a fair comparison, we follow the same setting as in previous work [45, 39, 35] using their official code and train or optimize each method under the same setup as ours  [^1] BPS, Physics-Net, PGAN and Two-Step are trained as they are learning-based methods; all other baselines are optimization-based; we train separate models for each noise level for both our method and other learning-based models. . Specifically, we compare our method with 3 traditional methods and 4 deep learning-based approaches: 1) BP [4]: A traditional non-iterative inversion algorithm. 2) 2-fold SOM [54]: A traditional iterative minimization scheme by using SVD decomposition. 3) Gs SOM [6]: A traditional subspace-based optimization method by decomposing the operator of Green’s function. 4) BPS [7, 45]: A CNN-based image translation method with an initial guess from the BP algorithm. 5) Physics-Net [26]: A CNN-based approach that incorporates physical phenomena during training. 6) Two-Step [51]: A CNN-based approach with two steps. 7) PGAN [39]: A CNN-based approach using a generative adversarial network. 8) Img-Interiors [27]: An implicit approach optimized by forward calculation. Following previous work [26], we evaluate the quantitative performance of our method using PSNR [44], SSIM [43], and Relative Root-Mean-Square Error (MSE[39].
@@ 5.2.1 Multiple Transmitter Evaluation
We begin by comparing our method against prior approaches under the multiple-transmitter setting, using both synthetic and real datasets for comprehensive evaluation. As shown in the upper part of Tab.˜1, our method achieves comparable or superior performance to the SOTA in most cases, demonstrating how our end-to-end training framework successfully leverage the data prior across diverse data domains. In addition, we present a qualitative comparison, as shown in Fig.˜4. Traditional methods such as BP[4], Gs SOM [6], and 2-fold SOM [6] are only capable of recovering the coarse shape of the scatterer. BPS [7, 45] produces sharp edges, but the reconstructed shapes are often inaccurate. PGAN [39] achieves accurate shape recovery, yet introduces noticeable background artifacts. Img-Interiors [27] can generate high-quality reconstructions, but occasionally fails due to local optima, as it is based on an iterative optimization process (see the last row). In contrast, our method produces accurate and clean reconstructions across all cases, demonstrating both visual fidelity and robustness.
@@ 5.2.2 Single Transmitter Evaluation
Figure 6: Qualitative comparison under single-transmitter setting on real-world IF dataset. The results are obtained with $N=1$ transmitter, without noise level. Colors represent the values of the relative permittivity.
Furthermore, we investigate a highly challenging and practically important setting that has been largely underexplored in previous work: performing EISP with a minimal number of transmitters. Specifically, we consider the most extreme case, using only a single transmitter. As shown in the lower part of Tab.˜1, our method significantly outperforms all previous approaches across all datasets and noise levels. This remarkable performance under such constrained conditions underscores the efficacy of our end-to-end training framework, which successfully encodes and leverages rich data priors to achieve state-of-the-art results across diverse domains. To better understand this phenomenon, we present qualitative comparisons in Fig.˜5 for synthetic data and Fig.˜6 for real-world IF [15]. Traditional methods such as BP[4], Gs SOM [6], and 2-fold SOM [6] produce only blurry reconstructions. Deep learning-based methods like BPS [7, 45], Physics-Net [26], and PGAN [39] tend to “hallucinate” the digit, resulting in wrong shape on the MNIST dataset. Img-Interiors [27] fails to capture the fundamental morphology of the scatterer, resulting in structurally inaccurate representations that deviate significantly from the ground truth. Among all the methods, only ours can still produce reasonably accurate reconstructions of the relative permittivity under such an extreme condition.
@@ 5.3 Ablation Study
Noise Robustness. To simulate real-world sensor noise and related perturbations, we evaluate the robustness of the models by adding noise to the scattered field. Moving beyond simple binary testing, we systematically assess the model’s performance across multiple noise levels ranging from 5% to 30% to examine its behavior in various noisy environments. The quantitative results presented in Tab.˜3 demonstrate that our model exhibits smooth and gradual performance degradation as the noise level increases, maintaining excellent reconstruction capability even under strong noise interference as high as 30%. Qualitative visualizations in Fig.˜7 show that most baseline methods exhibit noticeable artifacts or even complete failure under severe noise conditions, while our method remains robust and preserves the essential structure of the target.
@@
 Data MNIST (5%) MNIST (30%)
@@ 5.4 Reconstruction on 3D data
Setup and Metrics. Our method can be naturally extended to 3D scenarios. We use the same MLP architecture, with the input dimension extended to 3. Following previous work [27], we employ the Synthetic 3D MNIST [11] and extend to 3D ShapeNet [46] for training and testing. For evaluation, we adopt 3D versions of the MSE [39] and Intersection over Union (IoU) as our metrics. Further details on the datasets are provided in Appx.˜C.
Results. We evaluate our method and Img-Interiors under limited-transmitter settings. Quantitative results demonstrate the superiority of our approach: on 3D MNIST, our method achieves an MSE of 0.120 and IoU of 0.769 with $N=1$ transmitter, significantly outperforming Img-Interiors which obtains an MSE of 0.372 and IoU of 0.094 under the same conditions. With $N=6$ transmitters, our results further improve to MSE of 0.094 and IoU of 0.834. For the more complex 3D ShapeNet dataset under $N=1$ configuration, our method obtains an MSE of 0.064 and IoU of 0.762, showcasing its generalization capability to diverse 3D structures. Fig.˜1 provides a comprehensive comparison of reconstruction quality and runtime between the two methods for both $N=1$ and $N=6$ configurations. Fig.˜8 and Fig.˜9 provide visual comparisons of additional 3D reconstruction results on the 3D MNIST and 3D ShapeNet datasets. These results show that our method maintains robustness and generalizes effectively across geometrically complex 3D structures, representing significant progress towards practical applications.
@@ 6 Conclusion
Limitations. While our method effectively handles sparse transmitter settings, it struggles to reconstruct thin structures (see the rightmost block in Fig.˜9 for a typical failure case) and cannot accommodate varying receiver or transmitter locations. These limitations challenge deployment in real-world scenarios, where sensor layouts and environmental complexity vary. Addressing fine-structure reconstruction and flexible sensor configurations remains an important direction for future work.
Acknowledgments
This work was supported by NSFC-6247070125, Ningbo Grants 2025Z038 and 2025Z059, and the High Performance Computing Center at Eastern Institute of Technology and Ningbo Institute of Digital Twin.
@@ Appendix B Physical Model of EISP
In EISP, the scatterer is placed within the region of interest $\mathcal{D}$. The transmitters are placed around $\mathcal{D}$, emitting incident electromagnetic field ${E}^{\text{i}}$. The incident field then interacts with the scatterer, exciting the induced current $J$. The induced current can act as secondary radiation sources, emitting scattered field $E^{\text{s}}$. In fact, for a point in the scatterer, it cannot distinguish the incident field and the scattered field at this point. So, it interacts with the sum of the incident field and the scattered field, aka the total field $E^{\text{t}}$. The total field can be described by Lippmann-Schwinger equation [8] as follows:
@@
Since digital analysis only applies to discrete variables [28, 40], we discretize equations Eq.˜4, Eq.˜5, and Eq.˜6. The region of interest $\mathcal{D}$ is discretized into $M\times M$ square subunits, and we use the method of moment [33] to obtain the discrete scattered field $\mathbf{E}^{\text{s}}$ [^2] We use the bold letters to represent discrete variables.. The discrete version of Eq.˜4 is as follows:
@@ C.1 Datasets
The datasets utilized in this work are described in detail below. To enhance model robustness and training efficiency, datasets that share identical measurement settings are pooled to form consolidated training sets. We strictly follow the established rules in EISP [45, 39, 26] to set the dataset. We train and test our model on standard benchmarks used for EISP.
@@
1) Circular [27] is synthetically generated comprising images of cylinders with random relative radius, number, location, and permittivity between $1$ and $1.5$. $10k$ images are generated for training purposes, and $1.2k$ images for testing.
@@
2) MNIST [12] contains grayscale images of handwritten digits. Similar to previous settings [45, 57], we use them to synthesize scatterers with relative permittivity values between $2$ and $2.5$ according to their corresponding pixel values. The entire MNIST training set containing $60k$ images is used for training purposes, while $1.2k$ images from the MNIST test set are randomly selected for testing.
@@
Figure 10: Positions of transmitters and receivers for 2D data. (a) For Circular dataset and MNIST dataset, we set up $N=16$ transmitters and $N_{r}=32$ receivers. All transmitters and receivers are equally placed. For single-transmitter settings, the transmitter is positioned at the maximum x-coordinate at (3, 0). (b) For IF dataset, we set up $N=8$ or $18$ transmitters and $N_{r}=241$ receivers. The transmitters are equally placed (not shown in the figure), and the locations of receivers are determined by the transmitter. For all datasets, the green square represents the region of interest $\mathcal{D}$.
@@
Following previous work [7, 57] to generate the above two synthetic datasets, we set operating frequency $f=400$ MHz. The region of interest is a square with the size of $2m\times 2m$. We use $16$ transmitters and $32$ receivers equally placed on a circle with a radius of $3m$ ($N_{r}=32$ and $N=16$). The schematic diagram of the locations of the transmitters and receivers around the region of interest is shown in LABEL:fig:schematic_synthetic. The data are generated numerically using the method of moments [33] with a $224\times 224$ grid mesh to avoid inverse crime [9]. To simulate the noise in actual measurement, we add a $5\%$ level of noise to the scattered field $\mathbf{E}^{\text{s}}$ for regular setting.
@@
3) Real-world IF dataset [15] contains three different dielectric scenarios, namely FoamDielExt, FoamDielInt, and FoamTwinDiel. $N=8$ for FoamDielExt and FoamDielInt, $N=18$ for FoamTwinDiel, and $N_{r}=241$ for all the cases. The region of interest is a square with the size of $0.15m\times 0.15m$. Transmitters and receivers are placed on a circle with a radius of $1.67m$. The transmitters are placed equally, and the locations of receivers vary for each transmitter. There is no receiver at any position closer than $60^{\circ}$ from the transmitter, and 241 receivers are placed from $60^{\circ}$ to $300^{\circ}$ with a step of $1^{\circ}$ from the location of the transmitter. The schematic diagram of the locations of the transmitters and receivers around the region of interest is shown in LABEL:fig:schematic_IF. In the real measurement, there is only one transmitter at a fixed location, and the scatterer rotates to simulate the transmitter to be in different directions. There is a movable receiver sequentially measures the scattered field at 241 different locations. After the measurement, the scatterer rotates by a certain angle for next measurement. The angle is $45^{\circ}$ for FoamDielExt and FoamDielInt because $N_{r}=8$, and $20^{\circ}$ for FoamTwinDiel because $N_{r}=18$. As for operating frequency, all cases are measured under many different frequencies, and we take the frequency $f=5$ GHz. We evaluate these three scenarios for testing, and use the same settings to synthetically generate $10k$ images of cylinders with random number and location for training purposes.
@@
4) Synthetic 3D MNIST [11] contains 3D images of handwritten digits. 5) Synthetic 3D ShapeNet [46] contains 3D images of various shapes. We use these two datasets to synthesize scatterers with relative permittivity value of $2$. $N_{r}=160$ and $N=40$, and the operating frequency $f=400$ MHz. The region of interest is a cube with the size of $2m\times 2m\times 2m$. The transmitters and receivers are placed at a sphere with the radius of $3m$. For the positions of transmitters, the azimuthal angle ranges from $0^{\circ}$ to $315^{\circ}$ with the step of $45^{\circ}$, and the polar angle ranges from $30^{\circ}$ to $150^{\circ}$ with the step of $30^{\circ}$. For the positions of receivers, the azimuthal angle ranges from $0^{\circ}$ to $348.75^{\circ}$ with the step of $11.25^{\circ}$, and the polar angle ranges from $30^{\circ}$ to $150^{\circ}$ with the step of $30^{\circ}$. The schematic diagram of the locations of the transmitters and receivers around the region of interest is shown in Fig.˜11. The entire 3D MNIST dataset is used, including $5k$ images for training purposes and $1k$ images for testing. For 3D ShapeNet, we take $5k$ images from $5$ different categories for training purposes and $500$ images for testing.